OpenStudy (anonymous):

Tina lost 3 pounds on the first week of her diet. She gained a pound on the second week, and then lost 2 pounds a week during every week afterwards. She has been dieting for a total of 13 weeks. How many pounds has Tina lost in all? A. 26 pounds B. 20 pounds C. 24 pounds D. 22 pounds

OpenStudy (kittiwitti1):

lost 3 pounds: -3 gained 1 pound: +1 lost 2 pounds every week after: 2n (n is the consecutive weeks) Thirteen weeks minus two weeks is eleven weeks. So n is 11. Equation:\[-3+1-2n\rightarrow-3+1-2(11)=?\]

OpenStudy (kittiwitti1):

Are you still lost?

OpenStudy (anonymous):

yup

OpenStudy (kittiwitti1):

Where are you confused?

OpenStudy (anonymous):

i suck at math so all of it and i doing at home school so i don't really have a teacher to help me or teach me how to do any of this

OpenStudy (anonymous):

im gonna guess 24

OpenStudy (kittiwitti1):

1st week: lost 3 pounds. Lost means subtract, -3. 2nd week. gained 1 pound. Gain means add, +1 the other weeks: lost 2 pounds each. Lost means subtract. Since the first two weeks are of a different weight gain/loss, we subtract the first two weeks from the total 13 weeks and get 11 weeks, 13-2=11 weeks. So basically\[\text{week one}-3,\text{week two}+1\]\[\text{the remaining weeks}13-2=11\text{ weeks}\rightarrow-2n\text{with n being # of weeks}\]\[\text{since n is 11, }-2n=-2(11)\rightarrow\text{add it all up and }-3+1-2(11)\]

OpenStudy (kittiwitti1):

Maybe my explanation was confusing. But the equation you get to find the answer is -3+1-2(11)

OpenStudy (kittiwitti1):

\[-3+1-2(11)=-3+1-22=-25+1=-24\]You got it right, except it was negative. :]